What mass of NaCl is used to make 100.0 mL of a 0.0800 M NaCl solution?
Answer #1

We know that

Molariy= moles of NaCl/volume of solution in L

moles of NaCl = molarity*volume of solution in L = 0.0800*0.1= 8*10^-3 moles

mass of NaCl = moles*molar mass= 8*10^-3*58.44=0.467 gram NaCl